# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with multiplicative noise

### v1

@@ 4 Global Convergence of Sample Based Regularized Policy Gradient
   Sample a policy $\Sigma_{i}=\Sigma+V_{i}$, where $V_{i}$ is drawn uniformly at random over $\|V_{i}\|_{F}=r_{2}$.
@@
  Estimate: $\widehat{\nabla}_{\Sigma}=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{2}^{2}}f^{(l)}_{i}(K,\Sigma_{i})V_{i}$
@@
We define $f_{r_{1}}(K,\Sigma):=\mathbb{E}_{U\sim\mathbb{B}_{r_{1}}}[f(K+U,\Sigma)],$ and $f_{r_{2}}(K,\Sigma):=\mathbb{E}_{V\sim\mathbb{B}_{r_{2}}}[f(K,\Sigma+V)]$. where $\mathbb{B}_{r}$ denotes the uniform distribution over the points with norm $r$ (boundary of a sphere). The following lemma shows that the gradient of $f_{r_{1}}(K,\Sigma)$ and $f_{r_{2}}(K,\Sigma)$ can be estimated with an oracle for the function value.
@@
$\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)=\frac{n^{2}}{r_{2}^{2}}\mathbb{E}_{V\sim\mathbb{S}_{r_{2}}}[f(K,\Sigma+V)V].$
@@
The following lemmas show that $\nabla_{K}f(K,\Sigma)$, $\nabla_{\Sigma}f(K,\Sigma)$ and $S_{K,\Sigma}$ can be estimated with finite samples under small perturbation at any desired accuracy.
@@
Assume that (i) the distribution of the initial states implies that $\|x_{0}^{i}\leq L\|$ almost surely for any $i$. (ii) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$ for any $i$. If set
@@
with high probability (at least $1-\kappa_{1}$), where $\psi_{1}=Tr(\Sigma R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma|),\sigma_{1}=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2},R_{1}=\frac{2n}{r_{1}}f(K,\Sigma)+\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|},\sigma_{2}=(2\Gamma L^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2},R_{2}=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$.
@@
Given an arbitrary tolerance $\epsilon_{2}>0$ and probability $\kappa_{2}\in(0,1)$. Let $x_{t}^{i},u_{t}^{i}$ be $i$-th single path sampled using policy $(K,\Sigma+V_{i})\in\Omega$, where $\|V_{i}\|_{F}\leq r_{2}$, define
$\widehat{\nabla}_{\Sigma}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}.$
If set
$r_{2}$ $\leq\frac{\epsilon_{2}}{2h_{9}}$
$M$ $\geq\max\big\{\log\left(\frac{n}{\kappa_{2}}\right)\cdot\frac{4}{\epsilon_{2}^{2}}\left(2r_{2}^{2}+R_{2}\frac{\epsilon_{2}}{3}\right),\frac{6}{\epsilon_{2}^{2}}\log\left(\frac{n}{\kappa_{2}}\right)\left(3r_{2}^{2}+R_{2}^{\prime}\frac{\epsilon_{2}}{3}\right)\big\}$
$l$ $\geq\frac{\log\left(\frac{\epsilon_{2}r_{2}}{3n}\right){\log(\gamma)}-\log\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)+\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)|\psi_{2}|\right]}{\log(\gamma)}$
@@
with high probability (at least $1-\kappa_{2}$), where $\psi_{2}=\text{Tr}(\Sigma R)+\frac{\tau}{2}(n+2\log(2\pi)^{n}|\Sigma|)$ , $R_{2}=\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon_{2}}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$.
@@
$\leq\eta_{1}\frac{1}{\widehat{S}_{K,\Sigma}}\|\nabla_{K}f(K,\Sigma)-\widehat{\nabla}_{K}\|+\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}\Bigg|\frac{1}{S_{K,\Sigma}}-\frac{1}{\widehat{S}_{K,\Sigma}}\Bigg|$
@@
$\leq\eta_{2}\|\Sigma\|^{2}\|\nabla_{\Sigma}f(K,\Sigma)-\widehat{\nabla}_{\Sigma}\|$
@@ 5 Numerical Experiments
In this section we provide a 3 dimensional control and 1 dimensional state numerical experiment using SB-RPG (Algorithm 1), where all the parameters are unknown. Although SB-RPG does not require the exact values of the parameters, we need to specify the following parameters to simulate the system cost of $f(K,\Sigma)$. We set $A=0.7,B=(0.1,0.2,0.3),C=0.03,Q=0.5,\gamma=0.5,\tau=0.1,$
@@
The theoretical analysis provides rather conservative limits for the step size $\eta$, number of rollouts $M$, and rollout length $l$. To ensure practicality, we determined the constant step size, number of rollouts, rollout length, and exploration radius by performing a grid search over a set of reasonable values. In simulations, we obtained the baseline optimal cost $f(K^{*},\Sigma^{*})$ by solving the ARE in (2) to high accuracy ($e^{-5}$ accuracy) using value iteration.
@@ Acknowledgments
I first want to thank Dr. Xun Li for granting me the opportunity to join his research group and contribute to the meaningful work being done with his students. I would also like to extend my thanks to the Applied Math Department at The Hong Kong Polytechnic University for welcoming me during my two-weeks stay in Hong Kong.
@@ Appendix A Proofs in Section 2
$\quad+\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma P(B^{\top}B+D^{\top}D))u+\tau\log(\pi(u|x))+2\gamma APxBu\Bigr\}.$
@@
$=(2RK+\gamma P_{K}\nabla V_{K})\sum_{t=0}^{\infty}(\gamma V_{K})^{t}$
@@ Appendix B Proofs in Section 3
$C_{K^{\prime},\Sigma^{\prime}}(x_{0})-C_{K,\Sigma}(x_{0})$ $=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}-C_{K,\Sigma}(x_{0})$ (15)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})+C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}-C_{K}(x_{0})$ (16)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})\right]+\sum_{t=1}^{\infty}\gamma^{t}C_{K}(x^{\prime}_{t})\Biggl\}$ (17)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})+\gamma C_{K,\Sigma}(x^{\prime}_{t+1})-C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}$ (18)
$\triangleq\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x^{\prime}_{t},u_{t}^{\prime})\Biggl\}$ (19)
where $A_{K}(x,K^{\prime})$ is called ”advantage”, which can be viewed as the change in cost starting at state $x$ between one if choose $u_{t}=-K^{\prime}x$ only at current time and then $u_{t}=-Kx_{t}$ for all $t$ after the current time (i.e., $x^{2}(Q+K^{\prime\top}RK^{\prime})+C_{K}(x_{t+1})$) and one if $u_{t}=-Kx_{t}$ for all $t$ (i.e., $C_{K}(x)$)
We now want to find the $\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u^{\prime})\right]$
Expectation
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}$ (20)
$+\gamma\mathbb{E}\biggl\{P_{K}\left[(A+w^{x}C)x+(B+w^{u}D)u\right]^{2}\biggl\}-P_{K}x^{2}-q_{K,\Sigma}$ (21)
$=(Q+K^{\prime\top}RK^{\prime})x^{2}+Tr(\Sigma^{\prime}R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$ (22)
$+\gamma\left[P_{K}(A^{2}+C^{2}+K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-ABK^{\prime})x^{2}+Tr(\Sigma^{\prime}P_{K}(B^{\top}B+D^{\top}D))\right]$ (23)
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$ (24)
$+Tr(\Sigma^{\prime}(R+\gamma P_{K}(B^{\top}B+D^{\top}D)))$ (25)
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (26)
@@
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (28)
$=x^{2}[Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K^{\prime}}-(Q+K^{\top}RK+\gamma P_{K}V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (29)
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK+\gamma P_{K}(V_{K^{\prime}}-V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (30)
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK-2\gamma P_{K}AB(K^{\prime}-K)$ (31)
$\quad+\gamma P_{K}(K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (32)
@@
$\quad+\gamma P_{K}((K+K^{\prime}-K)^{\top}(B^{\top}B+D^{\top}D)(K+K^{\prime}-K)$ (33)
$\quad-\gamma P_{K}K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (34)
$=x^{2}[K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)$ (35)
$\quad+2(K^{\prime}-K)(R+\gamma P_{K}(B^{\top}B+D^{\top}D))K-\gamma P_{K}AB^{\top}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (36)
$=x^{2}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$ (37)
$+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (38)
@@
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $\geq-x^{2}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (39)
@@
$C_{K,\Sigma}(x_{0})-C_{K^{*},\Sigma^{*}}(x_{0})$ $=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{*}_{t},\pi^{*})\right]$
@@
$C_{K}(x_{0})-C_{K^{*}}(x_{0})$ $\geq C_{K}(x_{0})-C_{K^{\prime}}(x_{0})$ (41)
$=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{\prime}_{t},K^{\prime})\right]$ (42)
$=S_{K}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+h_{K}(\Sigma)-h_{K}(\Sigma^{\prime})$ (43)
$\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla_{K}C_{K,\Sigma}^{\top}(x_{0})\nabla_{K}C_{K,\Sigma}(x_{0})$ (44)

### v3

@@ 4 Global Convergence of Sample Based Regularized Policy Gradient
   Sample a policy $\Sigma_{i}=(L+V_{i})(L+V_{i})^{\top}$, where $V_{i}$ is drawn uniformly at random over $\|V_{i}\|_{F}=r_{2}$.
@@
  Estimate: $\widehat{\nabla}_{\Sigma}=\frac{L^{-1}}{2M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}f^{(l)}_{i}(K,\Sigma_{i})V_{i}$
@@
We define $f_{r_{1}}(K,\Sigma):=\mathbb{E}_{U\sim\mathbb{B}_{r_{1}}}[f(K+U,\Sigma)],$ and $f_{r_{2}}(K,L):=\mathbb{E}_{V\sim\mathbb{B}_{r_{2}}}[f(K,(L+V)(L+V)^{\top}]$. where $\mathbb{B}_{r}$ denotes the uniform distribution over the points with norm $r$ (boundary of a sphere). The following lemma shows that the gradient of $f_{r_{1}}(K,\Sigma)$ and $f_{r_{2}}(K,L)$ can be estimated with an oracle for the function value.
@@
$\nabla_{L}f_{r_{2}}(K,L)=\frac{n(n+1)}{2r_{2}^{2}}\mathbb{E}_{V\sim\mathbb{S}_{r_{2}}}[f(K,(L+V)(L+V)^{\top})V].$
@@
The following lemmas show that $\nabla_{K}f(K,\Sigma)$, $\nabla_{L}f(K,L)$ and $S_{K,\Sigma}$ can be estimated with finite samples under small perturbation at any desired accuracy.
@@
Assume that (i) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (ii) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$ for any $i$. If set
@@
with high probability (at least $1-\kappa_{1}$), where $\psi_{1}=Tr(\Sigma R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma|),\sigma_{1}=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2},R_{1}=\frac{2n}{r_{1}}f(K,\Sigma)+\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|},\sigma_{2}=(2\Gamma\bar{L}^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2},\\ R_{2}=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$.
@@
Given an arbitrary tolerance $\epsilon_{2}>0$ and probability $\kappa_{2}\in(0,1)$. Let $x_{t}^{i},u_{t}^{i}$ be $i$-th single path sampled using policy $(K,\Sigma_{i})\in\Omega$, where $\Sigma_{i}=(L+V_{i})(L+V_{i})^{\top}$, $L$ is choloky decomposition of $\Sigma$ such that $\Sigma=LL^{\top}$, $\|V_{i}\|_{F}\leq r_{2}$, define
$\widehat{\nabla}_{\Sigma}$ $:=\frac{1}{2}L^{-1}\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}$
Assume that (a) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (b) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}|x_{t})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right].$ If
$r_{2}$ $\leq\frac{\epsilon_{2}}{12h_{9}\|L^{-1}\|}$
$M$ $\geq\frac{2n(n+1)}{(\epsilon_{2}/3)^{2}}(\sigma_{Z}^{2}+\frac{R_{Z}{\epsilon_{2}/3}}{3})\log\big(\frac{n(n+1)/2+1}{1-\sqrt{1-\kappa}}\big)$
$l$ $\geq\frac{\log\frac{2r_{2}\epsilon_{2}}{6n(n+1)\|L^{-1}\|}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{\log\gamma}$
@@
with high probability (at least $1-\kappa_{2}$), where $R_{Z}=\max\{\frac{n(n+1)}{r_{2}}f(K,\Sigma)+\overline{\|\nabla_{L}f_{r_{2}}(K,L)\|}+\frac{\epsilon}{12\|L^{-1}\|},\\ \frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{\epsilon_{2}}{3\|L^{-1}\|}\},\sigma_{Z}=\max\{\frac{n(n+1)}{r_{2}}f(K,\Sigma),\big(\frac{n(n+1)}{r_{2}}\Gamma\bar{L}^{2}f(K,\Sigma)\big)^{2}\}$
@@
$\leq\eta_{1}\frac{1}{\widehat{S}_{K,\Sigma}}\|\nabla_{K}f(K,\Sigma)-\widehat{\nabla}_{K}\|+\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}\Bigg|\frac{1}{S_{K,\Sigma}}-\frac{1}{\widehat{S}_{K,\Sigma}}\Bigg|.$
@@
$\leq\eta_{2}\|\Sigma\|^{2}\|\nabla_{\Sigma}f(K,\Sigma)-\widehat{\nabla}_{\Sigma}\|.$
@@ 5 Numerical Experiments
In this section, we provide a numerical experiment using SB-RPG (Algorithm 1), where all the parameters are unknown. Although SB-RPG does not require the exact values of the parameters, we need to specify the following parameters to simulate the system cost of $f(K,\Sigma)$. We set $A=0.7,B=(0.1,0.2,0.3),C=0.03,Q=0.5,\gamma=0.5,\tau=0.1,$
@@
The theoretical analysis provides rather conservative limits for the step size $\eta$, number of rollouts $M$, and rollout length $l$. To ensure practicality, we determine the constant step size, number of rollouts, rollout length, and exploration radius by performing a grid search over a set of reasonable values. In simulations, we obtained the baseline optimal cost $f(K^{*},\Sigma^{*})$ by solving the ARE in (2) to high accuracy ($e^{-5}$ accuracy) using value iteration.
@@ Acknowledgments
The first two authors would like to thank the Department of Applied Mathematics at The Hong Kong Polytechnic University for the opportunity to join the research group and contribute to the meaningful work being carried out with its students.
@@ Appendix A Proofs in Section 2
$\quad+\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma P(B^{\top}B+D^{\top}D))u+\tau\log(\pi(u|x))+2\gamma APxBu\Bigr\}$
@@
$q_{K,\Sigma}=\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}.$
@@
$=(2RK+\gamma P_{K}\nabla V_{K})\sum_{t=0}^{\infty}(\gamma V_{K})^{t}.$
@@ Appendix B Proofs in Section 3
$C_{K^{\prime},\Sigma^{\prime}}(x_{0})-C_{K,\Sigma}(x_{0})$
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}-C_{K,\Sigma}(x_{0})$
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})+C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}-C_{K}(x_{0})$
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})\right]+\sum_{t=1}^{\infty}\gamma^{t}C_{K}(x^{\prime}_{t})\Biggl\}$
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})+\gamma C_{K,\Sigma}(x^{\prime}_{t+1})-C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}$
$\triangleq\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x^{\prime}_{t},u_{t}^{\prime})\Biggl\},$
where $A_{K}(x,K^{\prime})$ is called “advantage”, which can be viewed as the change in cost starting at state $x$ between one if choose $u_{t}=-K^{\prime}x$ only at current time and then $u_{t}=-Kx_{t}$ for all $t$ after the current time (i.e., $x^{2}(Q+K^{\prime\top}RK^{\prime})+C_{K}(x_{t+1})$) and one if $u_{t}=-Kx_{t}$ for all $t$ (i.e., $C_{K}(x)$). We now want to find $\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u^{\prime})\right]$.
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}$
$+\gamma\mathbb{E}\biggl\{P_{K}\left[(A+w^{x}C)x+(B+w^{u}D)u\right]^{2}\biggl\}-P_{K}x^{2}-q_{K,\Sigma}$
$=(Q+K^{\prime\top}RK^{\prime})x^{2}+Tr(\Sigma^{\prime}R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$
$+\gamma\left[P_{K}(A^{2}+C^{2}+K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-ABK^{\prime})x^{2}+Tr(\Sigma^{\prime}P_{K}(B^{\top}B+D^{\top}D))\right]$
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$
$+Tr(\Sigma^{\prime}(R+\gamma P_{K}(B^{\top}B+D^{\top}D)))$
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}}).$
@@
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
$=x^{2}[Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K^{\prime}}-(Q+K^{\top}RK+\gamma P_{K}V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK+\gamma P_{K}(V_{K^{\prime}}-V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK-2\gamma P_{K}AB(K^{\prime}-K)$
$\quad+\gamma P_{K}(K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
@@
$\quad+\gamma P_{K}((K+K^{\prime}-K)^{\top}(B^{\top}B+D^{\top}D)(K+K^{\prime}-K)$
$\quad-\gamma P_{K}K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
$=x^{2}[K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)$
$\quad+2(K^{\prime}-K)(R+\gamma P_{K}(B^{\top}B+D^{\top}D))K-\gamma P_{K}AB^{\top}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
$=x^{2}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
$+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
@@
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $\geq-x^{2}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
@@
$C_{K,\Sigma}(x_{0})-C_{K^{*},\Sigma^{*}}(x_{0})$
$=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{*}_{t},\pi^{*})\right]$
@@
$C_{K}(x_{0})-C_{K^{*}}(x_{0})$ $\geq C_{K}(x_{0})-C_{K^{\prime}}(x_{0})$
$=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{\prime}_{t},K^{\prime})\right]$
$=S_{K}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+h_{K}(\Sigma)-h_{K}(\Sigma^{\prime})$
$\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla_{K}C_{K,\Sigma}^{\top}(x_{0})\nabla_{K}C_{K,\Sigma}(x_{0})$